of the basketball is 14pi.What is the volume of the basketball


1
Expert's answer
2021-03-09T07:04:03-0500

Diameter of the basketball is = 14"\pi" cm

radius of the basketball is = "\frac{diameter}{2}" = "\frac{14\pi}{2}" = 7"\pi" = "7\times\frac{22}{7}" cm = 22cm


volume of the basketball is = "\frac{4}{3}\pi r^3" = "\frac{4}{3}\times\frac{22}{7}\times(22)^3" = 44,620.1905 cm3